# Disentangling Vortex-driven Dynamics in the Solar Atmosphere Using Information Theory

### v1

@@ Abstract
Solar tornadoes are believed to influence plasma dynamics and create conditions for heating, yet a direct quantitative link is lacking. Here, for the first time, we directly measure vortex-driven dynamics using information-theoretic diagnostics in a Bifrost simulation. By combining Shannon Entropy (SE) and Normalized Mutual Information (NMI), we track how vortices restructure plasma–magnetic interactions and channel energy into heat. The vortex flow originates in the upper photosphere and extends into the chromosphere and upper atmosphere. Relative to a control region dominated by shear flows and transient swirls, the coherent vortex shows stronger statistical interdependence between vorticity and other MHD variables. SE shows that vertical magnetic field entropy increases at the vortex onset as magnetic flux is redistributed, then decreases as the field becomes ordered. Magnetic shear and magnetic energy entropy peak during vortex development, reflecting current and energy build-up. In the upper atmosphere, low entropy in temperature and pressure alongside high entropy in density indicates ordered thermal and pressure fields but irregular mass distribution, a departure from ideal-gas behavior confirmed by weakened temperature–density coupling. NMI shows that temperature couples to different heating drivers with height: compression and vorticity (viscous heating, since vorticity traces velocity gradients) in the lower atmosphere, and vorticity plus magnetic shear in the upper atmosphere (viscous and current-driven heating). Although the simulation does not include explicit physical dissipation, hyperdiffusivity acts on sharp gradients and mimics these processes. Our results demonstrate that vortices drive multiscale coupling between flows and fields, locally shaping solar atmospheric dynamics and heating.
@@ 1 Introduction
The solar atmosphere supports a variety of vortices, from small-scale along intergranular lanes (Giagkiozis et al., 2018; Silva et al., 2020) to large vortical structures observed at the supergranular junctions (Requerey et al., 2017; Silva et al., 2018; Chian et al., 2019). The intergranular vortices may hit the upper atmosphere, displaying larger rotational signatures in the CaII and H core lines(Wedemeyer-Böhm et al., 2012; Dakanalis et al., 2021, 2022) known as chromospheric swirls. Wedemeyer-Böhm et al. (2012) traced observational signatures of vortices across the solar atmosphere showing that they form a structure called solar tornadoes. Silva et al. (2024b) identified that solar tornadoes are, in fact, composed of three co-spatial vortices: a classical flow vortex (the kinetic vortex Silva et al. (2021)), a magnetic vortex consisting of coherent twisted magnetic fields, and a Poynting flux vortex, characterized by swirling trajectories of energy flow.
@@
Numerical simulations have revealed a range of physical processes operating within vortex tubes. In high-plasma- layers, hydrodynamic flows can twist magnetic fields (Silva et al., 2021), generating magnetic shear and launching torsional waves (Fedun et al., 2011; Yadav et al., 2021). In contrast, in low- chromospheric regions, the dynamics can become magnetically dominated, with the magnetic field exerting control over plasma motions through magnetic tension and Lorentz forces (Silva et al., 2022, 2024a). Several studies have also highlighted the development of thin current sheets and localized heating at vortex boundaries, suggesting that magnetic shear and current dissipation may be central to vortex-related heating (Yadav et al., 2021; Silva et al., 2021). Vortices have also been linked to increased electromagnetic energy and plasma heating (Kuniyoshi et al., 2023, 2024; Silva et al., 2024a, b; Przybylski et al., 2025).
@@
While previous studies have demonstrated that vortices in the solar atmosphere often coincide spatially and temporally with magnetic field concentrations, plasma heating, shock and jet formation (e.g., Requerey et al., 2017; Tziotziou et al., 2023; Díaz-Castillo, S. M. et al., 2024), such associations remain largely qualitative. That is, they identify correlations based on co-location or apparent simultaneity, but they do not establish whether the vortex directly drives the observed dynamics. For instance, in the long-lived supergranular vortex studied by Requerey et al. (2017); Chian et al. (2022), the intensification and stabilization of a magnetic network element were observed to occur alongside vortex formation, but no formal causal analysis was performed. Consequently, these studies leave open key questions: Do vortices causally modulate heating or magnetic structuring, or are they simply co-evolving features responding to a common driver?
@@
In this study, we used IT tools to investigate a solar tornado captured in a 3D simulation of a coronal hole. Our goal is to go beyond visual associations and directly measure how key physical variables interact over time. IT tools, though only recently introduced to space plasma studies, have proven transformative in uncovering hidden structure and causality in nonlinear systems. Unlike traditional correlation-based diagnostics, IT methods can identify directionality and time-lagged interactions in turbulent multiscale environments. For example, transfer entropy has been used to show that solar flares may influence the timing of subsequent flaring events (Snelling et al., 2020), and to quantify how solar wind variability drives magnetospheric and radiation belt responses (Wing & Johnson, 2019). More recently, IT has revealed vortex–flow coupling in turbulent plasmas (Yatomi & Nakata, 2025), capturing causal pathways and scale-to-scale information transfer missed by conventional approaches. By computing both Shannon Entropy and normalized mutual information NMI , we are able to quantify, for the first time, the complexity of the dynamics in a simulated solar tornado and the strength of the coupling between vorticity and key MHD variables during the vortex flow lifecycle. We performed the same analysis in a region with a magnetic flux tube that did not support a solar tornado. This approach allows us to address several open questions: What mechanisms dominate the thermal evolution within large-scale solar vortices? How do magnetic shear and vorticity evolve during vortex formation and decay? Are heating processes localized and dynamic, or diffuse and passive? And, critically, can we separate correlation from causation in vortex-driven energy transfer? Through this lens, we demonstrate that magnetic structuring and current-related quantities are persistently and causally linked to plasma heating, particularly during the vortex phase. The enhanced statistical coupling in the vortex region, contrasted against a control case without a vortex, further emphasizes the dynamic and localized nature of these processes. The methods presented here provide a general framework for applying information theory to both simulations and observations of the solar atmosphere, offering new tools for diagnosing energy transport in highly nonlinear and data-limited environments.
We show that entropy increases and coupling patterns in the vortex region are consistent with magnetic complexity and dynamic interactions. The persistent NMI between temperature and magnetic shear during the vortex phase highlights current-driven heating as a dominant mechanism. Importantly, these coupling signatures vanish before and after the vortex event, confirming their dynamic origin. Our results establish IT diagnostics as powerful tools for probing causality and heating in nonlinear plasma environments and offer a path forward for future studies using simulations or limited observational datasets.
@@ 2.1 Numerical Model: Bifrost simulation
The simulation analysed in this study is dataset ch024031_by200bz005, produced using the Bifrost code (Gudiksen et al., 2011). The Bifrost code is a 3D radiative magnetohydrodynamics (MHD) model that captures the coupled dynamics of the solar convection zone, photosphere, chromosphere, transition region, and lower corona within a single, self-consistent framework. The analysed dataset consists of a vertically extended, low-density corona with a magnetically open topology. Thus, it captures key signatures of coronal holes, including reduced EUV emission and vertical magnetic connectivity. This simulation has also been used in earlier studies to examine vortex dynamics (Silva et al., 2024b; Finley, A. J. et al., 2022). The simulation domain spans Mm3, resolved with grid points, and extends from Mm below the solar surface (defined at optical depth unity, Mm) up to Mm into the corona. The lower boundary includes a steady horizontal magnetic inflow, injecting a uniform field of 5 G, superimposed on a net vertical field of 200 G. The simulation setup creates a strongly unipolar magnetic environment with open field lines. The upper boundary is open, and energy balance is handled through full radiative transfer in the lower atmosphere. This includes non-LTE scattering (Carlsson & Leenaarts, 2012), optically thin radiative losses, anisotropic thermal conduction, and ambipolar diffusion.
From the full simulation domain, we selected two sub-regions of Mm2 in the horizontal plane. One region hosts a solar tornado. This structure has been previously analyzed in detail by Silva et al. (2024b). The second region contains only a vertical magnetic flux tube and serves as a control region with no dominant vortical motion. To place the vortex analysis in spatial context, Figure 1 shows the full horizontal extent of the Bifrost simulation, colored by the vertical magnetic field component at a height of 2 Mm. The two selected subdomains are indicated by the black and red squares. The black square marks the region containing the large-scale, long-lived vortex, while the red square denotes the nearby control region. These contrasting subdomains allow us to assess the influence of vortex dynamics on the evolution of magnetic structure and plasma variability.
@@ 2.2 Dynamic Masking and Spatial Focus
To isolate the dynamics local to the kinetic vortex, we applied a spatial mask that is automatically tracking the evolving flux tube through height and time. The mask was defined as a circular region of fixed physical radius (in Mm), applied to each horizontal -slice of the vortex region. This mask served to limit all statistical calculations, namely SE, NMI, to the immediate neighborhood of the vortex core. The mask center was automatically determined for each height and time step as follows:
@@
To keep the mask aligned vertically, we set a limit on how much it could shift between height layers. If the new center moved more than two pixels from the one below it, we keep the old position instead. This helped to avoid jumping to nearby features or noise. We used the same rule when tracking the mask over time.
We ran this tracking method automatically across all time steps and heights, effectively building a 3D cylindrical region that follows the magnetic flux tube over time. Although vorticity was also computed, it was not used to define the tracking path. Magnetic flux concentrations appear before the vortex forms and remain after it fades, making them a more consistent reference. This approach also lets us compare regions with and without strong vortex activity to better isolate the effects of coherent rotation. An example of the selected region at 500 seconds and 3 Mm height is shown in Figure 2, with red arrows marking the direction of the velocity field and the location of the mask.
@@ 2.3 Information Theory Quantification
To understand how plasma properties vary and interact within the vortex, we applied IT tools to several physical variables, including the vertical magnetic field (), total magnetic field strength (), vertical velocity (), vorticity, temperature, density, and pressure. The first of these tools is Shannon entropy, which, in this context, quantifies the spatial variability or disorder of a physical quantity. A high entropy value indicates strong spatial fluctuations, pointing to a complex or possibly turbulent structure. In contrast, low entropy implies greater spatial uniformity, reflecting a more coherent or stable region.
Physically, high entropy in or suggests that the magnetic field is highly structured, possibly due to twisting, braiding, or the formation of current sheets. Low entropy in these quantities indicates an organized field, as seen in magnetically dominated regions or within stable flux tubes. For velocity and vorticity, high entropy reflects disordered or fine-structured flows. This can arise from turbulent motions or from overlapping wave modes, such as Alfvén or magnetoacoustic waves, which introduce spatial variability as they propagate through the magnetized plasma. When applied to thermodynamic variables such as temperature, density, and pressure, high entropy indicates significant spatial gradients and inhomogeneity—often associated with localized heating, shocks, or dynamic flows. Conversely, low entropy in these fields suggests more uniform conditions, typical of equilibrium or quasi-static regions.
In the mask region, each variable was normalized prior to histogram construction to ensure consistency across heights and time steps. To calculate entropy, we divide the variable range into bins and calculate the empirical probability as the fraction of grid cells that fall into the bin . The entropy was then computed in bits using a logarithmic base
@@
We tested both adaptive and fixed binning approaches. Adaptive binning adjusts to the local spread of the data, but it often over-resolved the histogram in regions with narrow distributions or limited sample sizes. This resulted in nearly uniform probability distributions and artificially high entropy values, which washed out spatial and temporal contrasts in the entropy maps. Fixed binning, on the contrary, preserved coherent structures and transitions across all variables and more clearly reflected known dynamic processes, such as vortex emergence and decay. For this reason, we adopt the fixed binning as the more interpretable and physically consistent method.
2.3.1 Coupling analysis using MI
MI is built on the same foundations as entropy, making it a powerful tool for uncovering relationships between variables. In both physics and information theory, entropy measures uncertainty or disorder. When two variables are related, knowing one reduces the uncertainty in the other. MI quantifies this reduction, not through assumptions about the type of relationship, but through the amount of shared information. This makes MI sensitive to both linear and nonlinear connections, even when other methods like correlation fail to capture them. Since it is rooted in entropy, a core idea in both thermodynamics and information theory, MI offers a broadly applicable and physically meaningful way to measure dependence between quantities.

### v2

# Characterizing Vortex-Driven Dynamics in the Solar Atmosphere Using Information Theory

@@ Abstract
Solar vortex regions show enhanced Poynting flux and favourable heating conditions, but how the vortices reorganize and influence their surroundings remains unclear Here we apply information-theoretic diagnostics to a Bifrost simulation to quantify the dynamics of a long-lived vortex. By combining Shannon Entropy and Normalized Mutual Information, we track how the vortex reshapes plasma-magnetic couplings and modifies local thermodynamics. The vortex originates in the upper photosphere and extends into the chromosphere, where it suppresses the background p-mode-like organisation seen in the neighbouring magnetic flux tube. Shannon Entropy analysis shows that magnetic complexity rises sharply as the vortex develops, which is consistent with the build-up of currents and stored energy. At the same time, temperature becomes more strongly linked to magnetic shear, pointing to heating associated with current dissipation. The way temperature responds to different heating processes also changes with height: in the photosphere and lower chromosphere, it follows local compressional and expansion motions, while in the upper atmosphere, it is influenced mainly by viscous and current-driven effects. During this phase, the usual temperature-density relationship weakens, indicating that the plasma departs from purely adiabatic behaviour. Applying the same diagnostics to a nearby non-vortical flux tube yields only weak, uniform couplings, which confirms that the enhanced links are vortex-driven. Together, these results demonstrate that a coherent solar vortex not only drives heating but also reconfigures the local atmosphere, replacing periodic pressure-driven behaviour with magnetically dominated dynamics.
@@ 1 Introduction
The solar atmosphere supports a variety of vortices, from small-scale along intergranular lanes (Giagkiozis_2018; Silva_2020; Yuan2023; Aljohani_2022; Shelyag_2011) to large vortical structures observed at the supergranular junctions (Requerey_2017; Silva_2018; Chian_2019). The intergranular vortices may hit the upper atmosphere, displaying larger rotational signatures in the CaII and H core lines (Wedemeyer2012; Dakanalis_2021; Dakanalis_2022) known as chromospheric swirls. Wedemeyer2012 traced observational signatures of vortices across the solar atmosphere, showing that they form structures often called solar tornadoes. Silva_2024a identified that solar tornadoes are, in fact, composed of three co-spatial vortices: a classical flow vortex (the kinetic vortex, Silva_2021), a magnetic vortex (Silva_2021) consisting of coherent twisted magnetic fields, and a Poynting flux vortex, characterized by swirling trajectories of energy flow (Silva_2024a).
@@
Numerical simulations have revealed a range of physical processes operating within vortex tubes. In high plasma- regions, plasma flows can twist magnetic fields (Silva_2021), generating magnetic shear and launching torsional waves (Fedun2011; Yadav_2021). In contrast, in low plasma- chromospheric regions, the dynamics can become magnetically dominated, with the magnetic field exerting control over plasma motions through magnetic tension and Lorentz forces (Silva_2022; Silva_2024b). Several studies have also highlighted the development of thin current sheets and localized heating at vortex boundaries, suggesting that magnetic shear and current dissipation may be central to vortex-related heating (Yadav_2021; Silva_2021; Silva_2024a; Kuniyoshi_2025). Vortices have also been linked to increased electromagnetic energy and plasma heating (Kuniyoshi_2023; Kuniyoshi_2024; Silva_2024b; Silva_2024a; przybylski2025).
@@
Previous studies have shown that vortices in the solar atmosphere often coincide spatially and temporally with magnetic field concentrations, plasma heating, shocks, and jet formation (e.g., Requerey_2017; Tziotziou_2023; Diaz_2024). However, the nature and extent of these associations remain largely qualitative, leaving a gap in quantitatively characterizing their physical relationships. That is, they identify correlations based on co-location or apparent simultaneity, but they do not establish whether the vortex directly drives the observed dynamics. For instance, in the long-lived supergranular vortex studied by Requerey_2017; chian_2023, the intensification and stabilization of a magnetic network element were observed to occur alongside vortex formation, but no formal causal analysis was performed. Consequently, these studies leave open a key question of whether vortices cause heating and magnetic structuring, or merely co-evolve with them.
@@
In this study, we use Information Theory (IT) methodology to investigate a solar tornado captured in a 3D simulation of a coronal hole. Our goal is to go beyond visual associations and directly measure how key physical variables interact over time. IT tools, though only recently introduced to space plasma studies, have proven transformative in uncovering hidden structure and causality in nonlinear systems. Unlike traditional correlation-based diagnostics, IT methods can identify directionality and time-lagged interactions in turbulent multiscale environments. For example, transfer entropy has been used to show that solar flares may influence the timing of subsequent flaring events (Snelling_2020), and to quantify how solar wind variability drives magnetospheric and radiation belt responses (Wing2019). More recently, IT has revealed vortex-flow coupling in turbulent plasmas (Yatomi2025), capturing causal pathways and scale-to-scale information transfer missed by conventional approaches.
By computing both Shannon Entropy (SE) and Normalized Mutual Information (NMI), we can quantify, for the first time, the degree of complexity of the plasma dynamics in the region where a solar tornado is present and also the strength of the coupling between vorticity and key MHD variables during the vortex flow lifecycle. To validate our results, we performed the same analysis focusing on a region with a magnetic flux tube without the presence of vortical flow. This multi-method approach allows us to address several open questions: (i) What mechanisms dominate the thermal evolution within large-scale solar vortices? (ii) How do magnetic shear and vorticity evolve during vortex formation and decay? (iii) Are heating processes localized and dynamic, or diffuse and passive? (iv) Critically, can we separate correlation from causation in vortex-driven energy transfer? Through information-theoric lens, we provide in this paper a novel way to address such critical questions. The methods presented here provide a general framework for applying information theory to both simulations and observations of the solar atmosphere, offering novel tools for diagnosing energy transport in highly nonlinear and data-limited environments. The paper is organized as follows. In the Methodology section, we introduce the numerical simulation used in our analysis, together with the mathematical framework of the IT methodology and the magnetic flux-tube tracking procedure. Our analysis of the dynamical complexity and coupling in the vortical region is presented in the Results section. Finally, in the Discussion and Conclusions, we further interpret our findings and summarize the main implications of this work.
@@ 2.1 Numerical Model: Bifrost simulation
The simulation employed in this study is the dataset ch024031_by200bz005, produced using the Bifrost code (Gudiksen_2011). The Bifrost code is a 3D resistive, radiative magnetohydrodynamics (MHD) model that captures the coupled dynamics of the solar convection zone, photosphere, chromosphere, transition region, and lower corona within a single, self-consistent framework. It employs sixth-order finite differences for spatial derivatives on a staggered mesh and third-order time integration with an adaptive time-step. The Bifrost model incorporates several key physical effects, including thermal conduction along magnetic field lines (Spitzer1957), Ohmic heating, viscous dissipation, and a comprehensive set of radiative heating and cooling processes. To manage numerical diffusion, Bifrost employs a split-diffusive scheme that separates the diffusive terms into a small, globally applied component and a localized hyperdiffusive component.
The analysed dataset consists of a vertically extended, low-density corona with a magnetically open topology. Thus, it captures key signatures of coronal holes, including reduced EUV emission and vertical magnetic connectivity. This simulation has also been used in earlier studies to examine vortex dynamics (Silva_2024a; Finley_2022). The simulation domain spans Mm3, resolved with grid points, and extends from Mm below the solar surface (defined at optical depth unity, Mm) up to Mm into the corona. The lower boundary includes a steady horizontal magnetic inflow, injecting a uniform field of 5 G, superimposed on a net vertical field of 200 G. The simulation setup creates a strongly unipolar magnetic environment with open field lines. The upper boundary is open, and the horizontal boundaries are periodic. The energy balance is handled through full radiative transfer in the lower atmosphere. This includes non-LTE scattering (Carlsson_2012), optically thin radiative losses, anisotropic thermal conduction, and ambipolar diffusion.
From the full simulation domain, we selected two sub-regions of Mm2 on the horizontal plane. One region hosts a solar tornado. This structure has been previously analyzed in detail by Silva_2024a. The second region contains only a vertical magnetic flux tube and serves as a control region, used as a reference to compare with the vortex case, as it exhibits no dominant vortical motion. To place the vortex analysis in spatial context, Figure 1 shows the full horizontal extent of the Bifrost simulation colored by the synthetic H core line. The spectra were computed accounting for departures from local thermodynamic equilibrium (non-LTE). To reduce the computational expense, we used SunnyNet (Chappell2022), which employs convolutional neural networks to approximate non-LTE hydrogen populations. The network was trained on two simulation snapshots generated with Multi3D (Leenaarts_2009). The two selected subdomains are indicated by the black and red squares. The black square marks the region containing the large-scale, long-lived vortex, while the red square denotes a nearby control region. In the close-up views, the -planes are colored by the vertical component of the magnetic field. The photospheric field tends to be highly concentrated at the footpoints of magnetic flux tubes, leading to a dark colour almost everywhere else. The flux tubes are indicated by the magnetic field lines in red. These contrasting subdomains allow us to assess the influence of vortex dynamics on the evolution of magnetic structures and plasma variability.
@@ 2.2 Dynamic Masking and Spatial Focus
To isolate the dynamics local to the kinetic vortex, we applied a spatial mask that automatically tracks the evolving flux tube through height and time. The mask was defined as a circular region of fixed physical radius (in Mm) applied to each horizontal -slice of the vortex region. This mask served to limit all statistical calculations, namely SE and NMI, to the immediate neighborhood of the vortex core. Although the vortex radius varies with height, the IT analysis required the regions to have the same size. This led to a compromise: an underestimation of the vortex region in the upper atmosphere and an overestimation near the photosphere. Nevertheless, the size of the analyzed region remains broadly consistent with the vortex extent for most of its vertical range.
The mask centre was automatically determined for each height and time step as follows:
@@
To keep the mask vertically aligned, we limited the displacement of its centre to within 1 Mm of its position in the adjacent height layer. At our spatial resolution (0.03125 Mm per pixel), this corresponds to a maximum shift of about 32 pixels. Candidate centres were first searched within a larger neighborhood (12 Mm radius) to avoid losing the structure, but any displacement exceeding the 1 Mm limit was rejected, and the previous centre was retained. This procedure prevents the mask from locking onto nearby structures or noise. We used the same displacement constraint for all time steps. This is justified by the short cadence (10 seconds) and the small horizontal drift of the vortex
We ran this tracking method automatically across all time steps and heights, effectively building a 3D cylindrical region that follows the magnetic flux tube over time. Magnetic flux concentrations appear before the vortex forms and remain after it fades away, making them a more consistent reference. This approach also lets us compare regions with and without strong vortex activity to better isolate the effects of coherent rotation. An example of the selected region at 500 seconds and 3 Mm height is shown in Figure 2, with red arrows marking the direction of the velocity field and the location of the mask.
@@ 2.3 Information Theory Quantification
To investigate the variability and statistical relationships among the plasma quantities within the vortex region, we applied IT tools to several variables, including the vertical magnetic field (), total magnetic field strength , vertical velocity (), vorticity, temperature (), density (), and pressure (). In the simulation domain, these quantities exhibit strong height dependence; therefore, within the masked vortex region, each variable was normalized to enable meaningful comparison across time and height. Because IT measures need discrete probability distributions, we converted each continuous field into a histogram within the mask.
To compute SE, the range of each variable was divided into bins, and the empirical probability, , was determined as the fraction of grid cells with values falling into bin , such that
The entropy was then calculated in bits (SE units) using,
@@
To determine the number of bins, , we tested both adaptive and fixed binning approaches. Adaptive binning adjusts to the local spread of the data, but it often over-resolves the histogram in regions with narrow distributions or limited sample sizes. This resulted in nearly uniform probability distributions and artificially high entropy values, which washed out spatial and temporal contrasts in the entropy maps. Fixed binning, on the contrary, preserved coherent structures and transitions across all variables and more clearly reflected known dynamic processes, such as vortex emergence and decay. For this reason, we adopt the fixed binning as the more interpretable and physically consistent method.
Mathematically, the SE measures the level of uncertainty associated with a probability distribution. If we compute the SE of a narrow Gaussian distribution where most values fall within a small range, the SE value will be low, as the outcome of a random sample can be predicted with high confidence. On the other hand, if the distribution is broader or flatter, as in a wide Gaussian, the probability is more evenly spread across a larger range of possible values, making it harder to predict the outcome, resulting in a higher entropy. In our analysis, SE of a 2D array quantifies the spatial variability or disorder of a physical quantity. A high entropy value indicates strong spatial fluctuations, related to a complex or possibly turbulent structure. In contrast, low entropy implies greater spatial uniformity, reflecting a more coherent or stable region. The lower panel of Figure 3 shows the SE computed at a height of 2.52 Mm for the plasma temperature within the magnetic flux tube that exhibits rotational plasma flow. For selected times, the upper panel of The lower panel of Figure 3 shows the SE computed at a height of 2.52 Mm for the plasma temperature within the magnetic flux tube that exhibits rotational plasma flow. For selected times, the upper panel of Figure 3 displays the magnitude of the temperature gradient, . Figure 3 illustrates that higher SE values (above 4) generally correspond to stronger and more irregular temperature distributions, as reflected in the corresponding temperature gradient maps. In contrast, lower SE values (equal to or below 2) are associated with more homogeneous temperature fields and correspondingly weaker gradients. The same reasoning applies to the other physical variables analyzed, where higher SE values indicate greater spatial complexity, higher gradients, and dynamical activity. For instance, the physical interpretation of high entropy in or suggests that the magnetic field is highly structured, possibly due to twisting, braiding, or the formation of current sheets. The low entropy in these quantities indicates an organized field, as seen in magnetically dominated regions or within stable flux tubes. For velocity and vorticity, high entropy reflects disordered or fine-structured flows. This can arise from turbulent motions or from overlapping wave modes, such as Alfvén or magnetoacoustic waves, which introduce spatial variability as they propagate through the magnetized plasma. When applied to thermodynamic variables such as density and pressure, high entropy indicates significant spatial gradients and inhomogeneity, often associated with localized shocks or dynamic flows. Conversely, low entropy in these fields suggests more uniform conditions, typical of equilibrium or quasi-static regions. The SE, thus, provides a quantitative diagnostic of the degree of structural organization and mixing within the magnetic environment, offering a more informative and objective measure for assessing how vortex dynamics influence the thermal and magnetic structuring of the plasma.
2.3.1 Quantifying statistical dependence with mutual information
MI is built upon the concept of entropy and provides a quantitative measure of the statistical dependence between variables. In both physics and information theory, entropy represents uncertainty or disorder, and MI quantifies how much knowledge of one variable reduces the uncertainty of another. In this sense, MI measures coupling strength, the degree to which two quantities share information. Unlike correlation, MI does not assume a specific functional form of the relationship, making it sensitive to both linear and nonlinear dependencies. Since MHD system are inherently nonlinear, traditional correlation-based measures may miss subtle, yet important dependencies. MI, however, is sensitive to the full spectrum of possible interactions as it stems from entropy, a concept central to both thermodynamics and information theory.
To compare how strongly different pairs of variables are connected, or how this connection changes across heights or times, MI is often normalized. This places all variable comparisons on a common scale, making it possible to assess the relative strength of statistical dependencies in a consistent manner. Several normalization schemes exist, such as dividing MI by the joint entropy or by the geometric mean of the individual entropies. In this study, we adopt the latter form,
which ensures symmetry between variables and yields values between 0 and 1. The geometric-mean normalization is a statistical choice introduced to enable comparison across variable pairs; it does not carry a direct physical meaning and may smooth out possible asymmetries in directional couplings. Once normalized, however, the resulting NMI values can still be interpreted in physical terms, as they express the fraction of shared information between two variables. For instance, an NMI value of 0.1 indicates that the two MHD variables share about 10% of their total information content.